# Tokens, the oft-overlooked appetizer: Large language models, the distributional hypothesis, and meaning

### v7

@@ 5 Post-Script
Finally, in our opinion, many recent findings lend support to the arguments we put forward here (first shared in preprint, December 2024). As discussed in Zimmerman (2025), Oh and Schuler (2024) Oh and Schuler (2024) and others provide convergent evidence regarding tokens as primitives – both as information receptacles and as mediators of distributional learning. For example, Gurnee and Tegmark (2024) report that “learned spatiotemporal representations are fairly robust to changes in prompting”, although “random distracting tokens or fully capitalizing entity names can degrade performance” Gurnee and Tegmark (2024), supporting our view that mapping a single concept across multiple tokenizations obscures conceptual unity for the model126126126By disaggregating information across tokens and hindering cross-context association of relevant distributional patterns. (see Figure 10). Similarly, Marjieh et al. (2025) describe conceptual entanglement arising from dual string–digit representations of numbers: “a representational tension in transformer models that must learn what a number is from text input” (a consequence of the aforementioned ‘flattening’). They demonstrate that “string bias can propagate into realistic settings, leading to incorrect answers” Marjieh et al. (2025). See also discussion of “evil numbers” Betley et al. (2025). McCoy et al. (2024) argue – consistent with our positions – that deep learning models should be understood through their objective functions,127127127They term this the ‘teleological approach’; we referred to it as ‘POSIWID teleology’ after Stafford Beer., that such models often struggle with atypical data, and that the conflation of frequency and truth, though powerful, carries substantial risks McCoy et al. (2024a). Lastly, Cui et al. (2024) identify a phase transition we interpret as potentially marking a shift between regime 1 (pattern-matching) and regime 2 (abstraction-based) learning. Their results suggest the interpretation that although tokens embody a commitment to linguistic primacy in cognition, models may be able to build larger abstractions atop that substrate, so that, practically, they may not be overly constrained by that choice Cui et al. (2024).
@@ 6 Limitations
The main limitations are human fallibility and their downstream functional implications (limited resources, limited concentration, limited knowledge, limited experience), the many approximations we relied on (e.g., in creating lists of words, in using a small corpus of fiction, in using automated parsers, etc.) which were then composed on each other, and the many simplifying assumptions (moving between ‘orthography’ and ‘language’; considering broad families of architecture as basically similar enough, e.g., generative LLMs like the GPTs and MLMs like RoBERTa;128128128We would like to explore their differences in future work. In some ways, generating the next token and guessing a masked token are similar tasks, but surely there ought to be less flexibility for an MLM’s predictions. What is the impact of that constraint on the model’s gnogeography? As span increases, e.g., as the model is asked to supply one, two, n tokens, does its gnogeography approach an LLMs? How much does constraint vary with syntactic position and role? How much does constraint vary, for any model, with corpus? For example, when disorganized speech is plausible (as in poetry or certain psychiatric contexts) or across registers or topics, how does the model’s gnogeography change? Presumably topics or formats that are more free-wheeling ought to engender more uniform probability distributions. moving between English examples and abstract considerations of language as a technology; using terms like ‘word-like’, ‘cognition’, and ‘meaning’ without providing precise definitions). We are happy to discuss why we made the choices we did with interested parties. See Appendix 42 for further discussion of research limitations.
7 Acknowledgements
@@
We used token to refer to the string representation of the symbolic portion of the model vocabularies, both before and after cleaning, the vector representation in the embedding, or both. For example, in the gnogeographic maps (e.g. Fig. 9), both the string (symbolic) and vector portions of the token are relevant. We tried to make it clear when each portion was being discussed, but noting here to help resolve potential ambiguity. Once the token is moving through the model, because of how the model works (notably the attention portions), it no longer represents the meaning of that token alone: it comes to contain significantly more information.
@@
We did some initial cleaning in order to make the vocabularies easier to read, and in order to consolidate them. In addition, we used tools such as tiktokenizer in order to view how input text could be parsed into tokens by a model Duong (2023).129129129This description of vocabulary treatment is essentially identical to one in the current draft of Cramer et al. 2024, in progress Cramer et al. (2024).
@@
Around May 2024, we further analyzed those vocabulary maps, and did some initial exploration with the vocabulary of the newly-released GPT-4o model, which we accessed as tiktoken’s “o200k_base”. There are six vocabularies available from tiktoken (gpt2, r50k_base, p50k_base, p50k_edit, cl100k_base, and o200k_base) OpenAI .  130130130Perhaps surprisingly, when we looked at four of the vocabulary maps produced via tiktoken, three (gpt2, p50k_base, and r50k_base) were identical. There could be differences in how the tokenizer breaks input sequences down into vocabulary components, or some other reason not visible in the vocabulary itself as to why these would be considered different vocabularies.
@@
For the initial straightforward comparisons among the tokens, we stripped whitespace and leading _ from tokens (used to indicate leading whitespace), but we did not strip other special symbols. Some vocabulary files included things like trailing w or leading ## (presumably indicating something about a word boundary).131131131E.g. ##rook, ##take, and ##washed are examples of tokens from funnel-transformer_small. For more detailed comparisons, we cleaned the files more aggressively, stripping from tokens anything that was not a space or in the Latin alphabet (upper- or lower-case).
@@
The HarperCollins 2019 word list (CSW19) HarperCollins (2019) is a list of words that are allowable in the English-language version of Scrabble. As far as we can tell, CSW19 is the most recent longest version of the list (words were removed in more recent versions), which makes it the most suitable for our purposes. We used CSW19 because it is highly-overseen, 132132132Funnily, overseen means the opposite of overlooked! meaning there are unlikely to be many mistakes and there should be relative consensus on the information in it, and because it provides consistent part-of-speech information.133133133We did find a few small mistakes or typos, which we corrected in our version of the list. Even so, it is not without caveats. We used the S2 list dwyl because it is more generous in its inclusion of words and contains lengths outside of the range 2-15.
@@
Applying the Kolmogorov-Smirnov test to the distributions of the proportions of parts of speech in the 193,905 POS tags for the CSW19 base words and the 86,889 cleaned tokens, we get a test statistic of 0.25 (0 means the CDFs are as close together as possible, 1 means they are maximally separated) and an unconvincing p-value of about 0.98, meaning we don’t have enough evidence to reject the null hypothesis that the distributions are different.
@@
Because there are rare words in the open classes (content words) that cause a heavy tail, we truncated the lists of tokens to see what the results would be in that case. See Fig. 24.
@@
Using Python, we retrieved spans of text around our exemplar tokens from a corpus of about 7,000 English-language works of fiction, most of which were in the science fiction or fantasy genre. We did minimal pre-processing of the text, which was from e-reader formats and fairly messy. For each exemplar token, we found every occurrence across the texts, excluded any that didn’t have sufficient context before and after the exemplar token, then shuffled them randomly. We used a BPE tokenizer to parse the text in order to find 100 tokens preceding and following the exemplar token.134134134As an interesting aside, Liu et al., 2019, the creators of RoBERTa, found “only slight differences” between a tokenizer with character-level resolution and one with byte-level resolution, even though the byte-level one should in general be more robust to OOV problems. They found the BPE encoding had “slightly worse end-task performance on some tasks”, but decided that the benefits of a “universal encoding scheme” outweighed the minor loss in performance. They called for “a more detailed comparison of these encodings” in future work – while we’re not comparing different tokenizations head-to-head, we hope this paper is spiritually in that vein, as we think one explanation for why a character-level vocabulary might have some benefits over a byte-level one could be found in our discussion of the token’s role as both vehicle for distributional patterns and semantic receptacle Liu et al. (2019).
@@
Sometimes identical strings appear multiple times as different tokens. This seems likely to be due to encoding. The collision could be happening at a number of levels, e.g. as the result of our cleaning processes, or it could be lower down, perhaps due to underlying choices in UTF-8 to accommodate relic systems.135135135If it is the latter, that could be a place where supplementary approximations of supradiegetic linguistic information could be helpful in facilitating the merger of tokens Zimmerman et al. (2024); tokenizers can have built-in merging functionality, depending on whether the algorithm is working at the byte or character level, so perhaps information about the sounds and orthography of the tokens involved could augment those processes. Depending on how the UTF-8 encoded character set is being handled, there could be scenarios where two or more different encodings would be be interpreted as the same character. Encoding is confusing and unwieldy; it seems like an area prone to human error (by the authors or others).
@@
Some of the common terms used to discuss LLMs can be ambiguous. For example, the meanings of “pre-training” and “training” can overlap, especially if fine-tuning also took place. Likewise, “character-level” can be defined in opposition to either smaller levels (the byte-level) or larger levels (subwords, words).136136136For example, compare usage in Liu et al., 2019 Liu et al. (2019) vs. Huang et al., 2023 Huang et al. (2023b). In the former case, a subword vocabulary could also be called a character-level vocabulary. We tried to interpret these terms appropriately when reading our sources, but we could have misunderstood.

### v8

@@ 5 Post-Script: what’s next
A growing body of work lends support to our arguments herein (first shared in preprint, December 2024). Several recent works support this view of tokens as foundational semantic and distributional primitives Oh and Schuler (2024). Gurnee and Tegmark (2024) show that learned spatiotemporal representations are robust to most prompting changes, yet performance degrades with random distracting tokens or atypical capitalizations, supporting the view that mapping a concept across multiple tokenizations obscures its unity for the model (see also Figure 10). Marjieh et al. (2025) show that conceptual entanglement arises from dual string-digit representations of numbers, leading to a “string bias” that can result in incorrect answers. Along these same lines, see discussion of ‘evil numbers’ in Betley et al. (2025). With respect to the role of the NTP objective function, McCoy et al. (2024a) likewise argue that deep learning models should be understood through their objective functions,127127127They term this the ‘teleological approach’; we referred to it as ‘POSIWID teleology’ after Stafford Beer., that such models often struggle with atypical data, and that the conflation of frequency and truth, though powerful, carries substantial risks.
Finally, Cui et al. (2024) identify a phase transition we interpret as marking a shift between pattern-matching (regime 1) and abstraction-based (regime 2) learning.128128128Consistent with our Discussion (Sec. 3), and further developed in Zimmerman (2025)., 129129129In conjunction, recall the earlier quote from Digutsch and Kosinski (2023): “GPT-3’s semantic activation is better predicted by similarity in words’ meaning (i.e., semantic similarity) rather than their co-occurrence in the language (i.e., associative similarity)”. Their results suggest that while LLMs embody a commitment to token-based linguistic primacy in cognition – with its attendant token-structure-based pitfalls – models may be able to build abstractions atop that substrate, so that, practically, they may not be overly constrained by that choice. This success is monumental: that LLMs can produce output that looks like ‘thinking’ through a cognition with (a form of) language baked into its core is revelatory for multiple fields (including philosophy, AI, and linguistics).
However, it also has profound implications for LLM ‘experience’. LLMs, with such curtailed bodies and access to the world as we know it, experience the world as and through text, which appears to them nihilogonously. As a result, all words are speech acts, and all actions are speech – an extreme, perhaps degenerative,130130130Maybe even perverse. instantiation of the idea that “all thoughts are thinkers” Levin (2024). Ultimately, this defines the gnogeographic ceiling of pure LLMs, and may continue to echo through their descendant iterations.131131131In other words, LLMs have – excitingly! – built abstract thought on a language-only foundation, but this success also traps them in a reality made entirely out of words, which fundamentally limits what they can be. This idea is further developed in Zimmerman et al. (2024)/ Zimmerman (2025).
@@ 6 Limitations
The main limitations are human fallibility and their downstream functional implications (limited resources, limited concentration, limited knowledge, limited experience), the many approximations we relied on (e.g., in creating lists of words, in using a small corpus of fiction, in using automated parsers, etc.) which were then composed on each other, and the many simplifying assumptions (moving between ‘orthography’ and ‘language’; considering broad families of architecture as basically similar enough, e.g., generative LLMs like the GPTs and MLMs like RoBERTa;132132132We would like to explore their differences in future work. In some ways, generating the next token and guessing a masked token are similar tasks, but surely there ought to be less flexibility for an MLM’s predictions. What is the impact of that constraint on the model’s gnogeography? As span increases, e.g., as the model is asked to supply one, two, n tokens, does its gnogeography approach an LLMs? How much does constraint vary with syntactic position and role? How much does constraint vary, for any model, with corpus? For example, when disorganized speech is plausible (as in poetry or certain psychiatric contexts) or across registers or topics, how does the model’s gnogeography change? Presumably topics or formats that are more free-wheeling ought to engender more uniform probability distributions. moving between English examples and abstract considerations of language as a technology; using terms like ‘word-like’, ‘cognition’, and ‘meaning’ without providing precise definitions). We are happy to discuss why we made the choices we did with interested parties. See Appendix 42 for further discussion of research limitations.
@@
We used token to refer to the string representation of the symbolic portion of the model vocabularies, both before and after cleaning, the vector representation in the embedding, or both. For example, in the gnogeographic maps (e.g. Figure 9), both the string (symbolic) and vector portions of the token are relevant. We tried to make it clear when each portion was being discussed, but noting here to help resolve potential ambiguity. Once the token is moving through the model, because of how the model works (notably the attention portions), it no longer represents the meaning of that token alone: it comes to contain significantly more information.
@@
We did some initial cleaning in order to make the vocabularies easier to read, and in order to consolidate them. In addition, we used tools such as tiktokenizer in order to view how input text could be parsed into tokens by a model Duong (2023).133133133This description of vocabulary treatment is essentially identical to one in the current draft of Cramer et al. 2024, in progress Cramer et al. (2025).
@@
Around May 2024, we further analyzed those vocabulary maps, and did some initial exploration with the vocabulary of the newly-released GPT-4o model, which we accessed as tiktoken’s “o200k_base”. There are six vocabularies available from tiktoken (gpt2, r50k_base, p50k_base, p50k_edit, cl100k_base, and o200k_base) OpenAI .  134134134Perhaps surprisingly, when we looked at four of the vocabulary maps produced via tiktoken, three (gpt2, p50k_base, and r50k_base) were identical. There could be differences in how the tokenizer breaks input sequences down into vocabulary components, or some other reason not visible in the vocabulary itself as to why these would be considered different vocabularies.
@@
For the initial straightforward comparisons among the tokens, we stripped whitespace and leading _ from tokens (used to indicate leading whitespace), but we did not strip other special symbols. Some vocabulary files included things like trailing w or leading ## (presumably indicating something about a word boundary).135135135E.g. ##rook, ##take, and ##washed are examples of tokens from funnel-transformer_small. For more detailed comparisons, we cleaned the files more aggressively, stripping from tokens anything that was not a space or in the Latin alphabet (upper- or lower-case).
@@
The HarperCollins 2019 word list (CSW19) HarperCollins (2019) is a list of words that are allowable in the English-language version of Scrabble. As far as we can tell, CSW19 is the most recent longest version of the list (words were removed in more recent versions), which makes it the most suitable for our purposes. We used CSW19 because it is highly-overseen, 136136136Funnily, overseen means the opposite of overlooked! meaning there are unlikely to be many mistakes and there should be relative consensus on the information in it, and because it provides consistent part-of-speech information.137137137We did find a few small mistakes or typos, which we corrected in our version of the list. Even so, it is not without caveats. We used the S2 list dwyl because it is more generous in its inclusion of words and contains lengths outside of the range 2-15.
@@
Because there are rare words in the open classes (content words) that cause a heavy tail, we truncated the lists of tokens to see what the results would be in that case. See Figure 24.
@@
Using Python, we retrieved spans of text around our exemplar tokens from a corpus of about 7,000 English-language works of fiction, most of which were in the science fiction or fantasy genre. We did minimal pre-processing of the text, which was from e-reader formats and fairly messy. For each exemplar token, we found every occurrence across the texts, excluded any that didn’t have sufficient context before and after the exemplar token, then shuffled them randomly. We used a BPE tokenizer to parse the text in order to find 100 tokens preceding and following the exemplar token.138138138As an interesting aside, Liu et al., 2019, the creators of RoBERTa, found “only slight differences” between a tokenizer with character-level resolution and one with byte-level resolution, even though the byte-level one should in general be more robust to OOV problems. They found the BPE encoding had “slightly worse end-task performance on some tasks”, but decided that the benefits of a “universal encoding scheme” outweighed the minor loss in performance. They called for “a more detailed comparison of these encodings” in future work – while we’re not comparing different tokenizations head-to-head, we hope this paper is spiritually in that vein, as we think one explanation for why a character-level vocabulary might have some benefits over a byte-level one could be found in our discussion of the token’s role as both vehicle for distributional patterns and semantic receptacle Liu et al. (2019).
@@
Sometimes identical strings appear multiple times as different tokens. This seems likely to be due to encoding. The collision could be happening at a number of levels, e.g. as the result of our cleaning processes, or it could be lower down, perhaps due to underlying choices in UTF-8 to accommodate relic systems.139139139If it is the latter, that could be a place where supplementary approximations of supradiegetic linguistic information could be helpful in facilitating the merger of tokens Zimmerman et al. (2024); tokenizers can have built-in merging functionality, depending on whether the algorithm is working at the byte or character level, so perhaps information about the sounds and orthography of the tokens involved could augment those processes. Depending on how the UTF-8 encoded character set is being handled, there could be scenarios where two or more different encodings would be be interpreted as the same character. Encoding is confusing and unwieldy; it seems like an area prone to human error (by the authors or others).
@@
Some of the common terms used to discuss LLMs can be ambiguous. For example, the meanings of “pre-training” and “training” can overlap, especially if fine-tuning also took place. Likewise, “character-level” can be defined in opposition to either smaller levels (the byte-level) or larger levels (subwords, words).140140140For example, compare usage in Liu et al., 2019 Liu et al. (2019) vs. Huang et al., 2023 Huang et al. (2023b). In the former case, a subword vocabulary could also be called a character-level vocabulary. We tried to interpret these terms appropriately when reading our sources, but we could have misunderstood.